What is a torsional pendulum ? Derive the
expression for the time period of a torsional
pendulum. State one use of such a
pendulum.
1
Expert's answer
2019-06-06T10:02:55-0400

A torsional pendulum consists of a disk (or some other rigid body) suspended by a light wire or spring, which is then twisted and released, resulting in an oscillatory motion. 

The oscillatory motion is caused by a restoring torque τ which is proportional to the angular displacement θ:


"\tau =-k\theta"

where k>0  is the torque constant of the wire or spring. The minus sign shows that the restoring torque acts in the opposite direction to increasing angular displacement. This equation is a torsional equivalent to Hooke's law. The net torque is equal to the moment of inertia of the disk I times the angular acceleration:

"I\frac{{{d}^{2}}\theta }{d{{t}^{2}}}=\tau"


Combining the previous two equations, we obtain


"I\frac{{{d}^{2}}\theta }{d{{t}^{2}}}=-k\theta"

or


"\frac{{{d}^{2}}\theta }{d{{t}^{2}}}+\frac{k}{I}\theta =0"

This equation is a simple harmonic equation. Therefore, we can immediately write a standard solution.


"\theta ={{\theta }_{0}}\cos \left( \omega t+\varphi \right)"

where "{{\theta }_{0}}" is the maximum angle, "\varphi" is the phase, "\omega =\sqrt{\frac{k}{I}}" is the angular velocity. The period of oscillation of the pendulum is


"T=\frac{2\pi }{\omega }=2\pi \sqrt{\frac{I}{k}}"

Torsion pendulum is used in mechanical wristwatch. The balance wheel in a mechanical wristwatch is a torsion pendulum in which the restoring torque is provided by a coiled spring.